Subject: APO 3.5m users committee minutes, Nov 2011

From: Suzanne Hawley

Submitted: Fri, 11 Nov 2011 15:22:12 -0800 (PST)

Message number: 1251 (previous: 1250, next: 1252 up: Index)

 		APO 3.5-m Users Committee Phonecon, 11/07/11

Attending:  Suzanne Hawley, Rich Kron (for Al Harper), Remy Indebetouw, 
John Bally, Michael Strauss, Scott Anderson, Sean Moran, and Bruce 

Absent:  Mark Klaene, Bill Ketzeback, Jon Holtzman


User feedback and comments from institutional representatives:

o Washington (Scott Anderson) - Scott had nothing to report.

o Colorado (John Bally) - John had nothing to report.

o Johns Hopkins (Sean Moran) - Sean reported that one of his JHU users saw 
fringing in SPIcam data, and that there appears not to be any user 
information on our website about it.  Suzanne said that SPIcam is 
susceptible to fringing especially in i and z, and that we usually let 
users deal with it using standard IRAF routines.  Bill Ketzeback will
update the documentation with information about what is to be expected 
from SPICam.  We anticipate that GRIZLI will not have significant 
fringing, as its detectors are deep-depletion devices.

o New Mexico State (Jon Holtzman) - No report.

o Princeton (Michael Strauss) - Michael said that everything was going OK. 
Suzanne asked if the Princeton ToO program was getting good data, and 
Michael said yes.

o Virginia (Remy Indebetouw) - Remy had nothing to report.

o Chicago (Rich Kron) - Rich asked a question from one of the UC users 
regarding the status of the echelle front end.  We took this to mean the 
slit camera, which has been especially cranky in recent months.  Suzanne 
said that a new camera and computer are on order to replace this ancient, 
unmaintainable system.  In the meantime, some software and network patches 
were implemented that seem to have gotten the camera to a state of 
reasonable reliability.

** Note added 11/11/11: The echelle slitviewer camera had a shutter 
problem this week which was fixed by ordering and installing a new 
shutter.  It should now be operational again.


Discussion of telescope/instruments report:

Suzanne said that we've enjoyed a remarkable year of good observing 
weather.  Whereas we typically lose ~35% of all observing time to weather, 
last year it was only 24%.  The telescope has been working well, and the 
seasonal behind-the-scenes adjustments are underway to tune it for 
cold-weather operation.  A new B400 grating arrived at JHU and is being 
tested; we anticipate putting it into service in early December, and users 
will be warned to be on the lookout for differences in performance between 
the new and old gratings.  We are also continuing the research for a 
possible new grating for the echelle, and Steve Smee and Robert Barkhouser 
plan a trip to APO to look at mounting design issues.  Most of the staff 
engineering work in the past month was on NIC-FPS.  It is still having 
readout problems, but we currently are letting users observe with the 
instrument on a shared-risk basis.  We also have a new display routine 
that shows all the frames taken in a Fowler read so that read failures can 
be detected quickly.  Suzanne is investigating establishing a long-term 
maintenance plan for NIC-FPS, possibly involving the UVa instrument group. 
Lastly, Agile was recently put back into use after a hiatus, and problems 
with its fiber communications were tracked down and fixed.

Mark Klaene's detailed report for the reporting period follows:


 	      3.5-m Telescope and Instruments Highlights
 	                 9/28/11 through 11/3/11

0) Overview

We completed the GRIZLI (GRIZ Lambda Imager, previously known as PISCO) 
preliminary design review in October.  On site reviewers included Fred 
Hearty and John Wilson from UVa and Jon Holtzman from NMSU.  Members from 
other institutions are reviewing the documentation off line. An NMSU class 
trip took place in October and a large UVa class trip is scheduled for 
this month. Reminder, the Alamo/El Paso Shuttle has changed owners. 
Rates and schedule have changed but their contact information remains the 
same. A dry and relatively clear fall is upon us with more of the same 
forecasted for the winter.

1) Telescope

The telescope has been operational with generally only minor problems 
reported. The moths made one last stand before the cold set in, causing 
some minor issues.  A temporary loss of the 1 pps signal also required a 
controller reset that affected slews and tracking. As the temperature 
drops and the telescope equilibrates to a lower structural temperature, 
minor adjustments may be required to preloads and servo parameters.

2) Instruments

DIS replacement Blue 400 grating has been ordered and is expected here 
around early December. Isolated problems with the DIS robotics have 
occurred and caused some loss time. This is not uncommon as the robotics 
electronics have a long history of intermittent problems.

TSpec and SPIcam operated with no new issues. Echelle is operational. An 
MOU with JHU for the replacement grating order is in work. The ECAM 
(echelle slitviewer) problem has been improved significantly with some 
software modifications and network upgrades. The replacement camera order 
was delayed in purchasing and will not arrive until next year.

NIC-FPS continued to be the issue of the month.  Staff continued 
troubleshooting the instrument while learning the intricacies of the 
detector electronics. A bad fiber was also found between the ICC and the 
Leach controller.  The icing of the bulkhead connector has been 
significantly reduced and possibly eliminated by a modification to the 
warm side of the connector.  The image corruption issue is still with us 
but seems to be highly variable.  There appears to be no logic as to when 
and how the images fail.  Consequently we have returned the instrument to 
shared-risk observing status to help gather data on the frequency of the 
failures while collecting what we believe to be usable science data.

Agile lost the timing signal from the ICC.  The fiber to the computer room 
and the fiber to TTL converter were replaced, and it appears to be 

GFP software development is continuing with a developer on site this week.


GRIZLI update:

Suzanne gave an update on the newly named GRIZLI (thanks to Rene for the 
name which stands for GRIZ Lambda Imager, the instrument which was known 
as PISCO in its Magellan version) which is a new multi-band imager being 
built by Chris Stubbs and Tony Stark at Harvard/CfA.  A PDR for the 
instrument was held at APO a couple of weeks ago, and the results were 
favorable.  We expect a delivery at APO sometime in 2012 if all goes well, 
and in the meantime we are exploring scattered light issues in the 
telescope that may or may not be a problem for this wide-field camera. 
Bruce noted that given the large(r) FOV plus 4-color multiplexing, GRIZLI 
can be thought of as a significant fraction of the ODI instrument being 
built for WIYN.


Target of Opportunity update for Q1 scheduling:

Suzanne referred to the ToO program we are running this quarter for a 
Princeton astronomer, who has access to PTF data.  We will continue the 
ToO opportunities into next year for anyone interested, and Suzanne said 
that ToO preemptions will be limited to approximately 3 per quarter per 
user (1 per month).  For Q1 scheduling, if a user plans to ask for 3 ToOs, 
that user should have 3 dark halfs allocated by the institution to be used 
for payback.  Also, all should be aware that APO policy is that the PI of 
the pre-empted program is invited to collaborate and be an author on any 
publication that results from data taken during the ToO.  Other details on 
the ToO policies and procedures are being reviewed and updated; Suzanne 
will bring these forward for review soon.


NIR imaging - NIC-FPS and NIIS status:

NIC-FPS was discussed previously in the meeting. Suzanne mentioned that 
NIIS (Near-IR Imaging Spectrograph - which doesn't yet have a 
spectrograph, like NIC-FPS doesn't have an FPS) is a wide-field NIR imager 
being built by Michael Pierce at UWyoming.  We are planning for it to come 
to APO as a visiting instrument for a week in March 2012.  As we get 
closer to the date, we will determine if we can try to take commissioning 
observations on a shared-risk basis.  Users committee members have the 
instrument details, so users should ask them if they would like to try to 
get some early data.  It is unclear if we can ever use NIIS as a facility 
instrument, mainly because of its size -- it would displace two other 
instruments if we were to have it available routinely.  One possibility 
would be to schedule its future use in campaign mode.


References needed for annual report:

Suzanne said that Rene Walterbos has in the past graciously compiled the 
annual list of 3.5-m publications, theses, etc., for the annual Board 
meeting.  Now that Rene is the ARC Board Chair, and heavily involved in 
AS3 and other ARC futures developments, we need to assemble the lists of 
telescope publications ourselves.  We don't need these for next week's 
Board meeting, but we will need them by the end of the year.

===> ACTION:  User committee members to assemble a listing of all 3.5-m 
telescope-related publications, theses, etc., published by users at their 
institutions from the past year, October 2010 through September 2011, and 
send list to Suzanne by 12 December.



[open from previous meeting]:

===> ACTION:  Users Committee members to provide Mark and Suzanne with 
institutional science highlight for the year, by 17 October, 1 slide or 
figures plus text.
      STATUS:  Discussed at this meeting, closed.

[new actions from this meeting]:

===> ACTION: User committee members to assemble a listing of all 3.5-m 
telescope-related publications, theses, etc., published by users at their 
institution from the past year, October 2010 through September 2011, and 
send list to Suzanne by 12 December.


Next meeting:  The next Users Committee phonecon will be on Monday 12 
December at 8:30 AM Pacific Time.  The agenda and other materials will be 
sent to the committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1251 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to